### v3

@@ Introduction
[…]
@@
micro avg
(support: 8598)
recall 0.766 0.768 0.788 0.783 0.859 0.856 0.868 0.859 0.874 0.879
@@
macro avg
(support: 8598)
recall 0.728 0.729 0.747 0.748 0.834 0.830 0.844 0.832 0.853 0.855
@@
weighted avg
(support: 8598)
recall 0.766 0.768 0.788 0.783 0.859 0.856 0.868 0.859 0.874 0.879
@@
MISC
(support: 1029)
recall 0.504 0.500 0.508 0.548 0.690 0.684 0.706 0.673 0.727 0.718
@@
micro avg
(support: 8598)
recall 0.874 0.879 0.843 0.848 0.854 0.849 0.852 0.854 0.870 0.869 0.877 0.876 0.879 0.875 0.870 0.874 0.876 0.876 0.866 0.869 0.872 0.867 0.862 0.871 0.870 0.864 0.860 0.857 0.873 0.872
@@
macro avg
(support: 8598)
recall 0.853 0.855 0.813 0.818 0.826 0.815 0.821 0.826 0.840 0.840 0.851 0.850 0.855 0.850 0.842 0.847 0.851 0.847 0.842 0.841 0.846 0.839 0.836 0.844 0.843 0.835 0.830 0.825 0.848 0.844
@@
weighted avg
(support: 8598)
recall 0.874 0.879 0.843 0.848 0.854 0.849 0.852 0.854 0.870 0.869 0.877 0.876 0.879 0.875 0.870 0.874 0.876 0.876 0.866 0.869 0.872 0.867 0.862 0.871 0.870 0.864 0.860 0.857 0.873 0.872
@@
MISC
(support: 1029)
recall 0.727 0.718 0.640 0.648 0.666 0.624 0.645 0.667 0.668 0.674 0.703 0.705 0.721 0.706 0.684 0.697 0.709 0.679 0.698 0.683 0.693 0.679 0.680 0.686 0.694 0.668 0.659 0.643 0.702 0.686

### v4

@@ Introduction
$P(\text{entity tag}|\text{token})=\text{softmax}(W_{c}h+b_{c})$
@@
micro avg (support: 8598) recall 0.766 0.768 0.788 0.783 0.859 0.856 0.868 0.859 0.874 0.879
@@
macro avg (support: 8598) recall 0.728 0.729 0.747 0.748 0.834 0.830 0.844 0.832 0.853 0.855
@@
weighted avg (support: 8598) recall 0.766 0.768 0.788 0.783 0.859 0.856 0.868 0.859 0.874 0.879
@@
MISC (support: 1029) recall 0.504 0.500 0.508 0.548 0.690 0.684 0.706 0.673 0.727 0.718
@@
micro avg (support: 8598) recall 0.874 0.879 0.843 0.848 0.854 0.849 0.852 0.854 0.870 0.869 0.877 0.876 0.879 0.875 0.870 0.874 0.876 0.876 0.866 0.869 0.872 0.867 0.862 0.871 0.870 0.864 0.860 0.857 0.873 0.872
@@
macro avg (support: 8598) recall 0.853 0.855 0.813 0.818 0.826 0.815 0.821 0.826 0.840 0.840 0.851 0.850 0.855 0.850 0.842 0.847 0.851 0.847 0.842 0.841 0.846 0.839 0.836 0.844 0.843 0.835 0.830 0.825 0.848 0.844
@@
weighted avg (support: 8598) recall 0.874 0.879 0.843 0.848 0.854 0.849 0.852 0.854 0.870 0.869 0.877 0.876 0.879 0.875 0.870 0.874 0.876 0.876 0.866 0.869 0.872 0.867 0.862 0.871 0.870 0.864 0.860 0.857 0.873 0.872
@@
MISC (support: 1029) recall 0.727 0.718 0.640 0.648 0.666 0.624 0.645 0.667 0.668 0.674 0.703 0.705 0.721 0.706 0.684 0.697 0.709 0.679 0.698 0.683 0.693 0.679 0.680 0.686 0.694 0.668 0.659 0.643 0.702 0.686